OpenStudy (anonymous):

Will medal, solve the equation 4sin theta - sqrt 3 = 2sin theta. express in radian measure.

OpenStudy (danjs):

\[4\sin(\theta) - \sqrt{3}= 2 \sin(\theta)\] that righ

OpenStudy (danjs):

Or is there a parenthesis around (theta - sqrt 3) \[4\sin(\theta-\sqrt{3}) = 2\sin(\theta)\]

OpenStudy (anonymous):

first one is correct, no parenthesis @danjs